A researcher claims that the monthly consumption of coffee per person is more than 19 cups. In a sample of 60 randomly selected people, the mean monthly consumption was 20. The standard deviation of the sample was 4 cups. Find the p-value of the test. On the basis of the p-value, is the researchers claim valid at 0.01 level of significance?


1
Expert's answer
2021-09-29T15:48:56-0400

"H_0: \mu≤ 19 \\

H_1: \mu>19 \\

\bar{x} = 20 \\

n=60 \\

s=4"

Test-statistic:

"t = \frac{\bar{x} -\mu}{ s / \sqrt{n}} \\

t = \frac{20-19}{4 / \sqrt{60}} =1.937"

The EXCEL formula to find the p-value for a two-tailed t-test and df=60-1=59 is

=tdist(1.937, 59, 1)

p-value = 0.0287

p>0.01

At significance level of 1%, the researcher can NOT claim that the monthly consumption of coffee per person is more than 19 cups.